A noodle machine in Spumoni’s spaghetti factory makes about 5 percent defective noodles even when properly adjusted. The noodles are then packed in crates containing 1900 noodles each. A crate is examined and found to contain 115 defective noodles. What is the approximate probability of finding at least this many defective noodles if the machine is properly adjusted?
Answer #1

ANSWER::

n = 1900

p = 0.05

q = 0.95

Mean = np = 95

Standard deviation = = 9.5

P(X < A) = P(Z < (A - mean)/standard deviation)

P(at leas 115 defective noodles) = P(X > 114.5)

= 1 - P(X < 114.5)

= 1 - P(Z < (114.5 - 95)/9.5)

= 1 - P(Z < 2.05)

= 1 - 0.9798

= 0.0202

NOTE:: I HOPE YOUR HAPPY WITH MY ANSWER....***PLEASE SUPPORT ME WITH YOUR RATING...

***PLEASE GIVE ME "LIKE"...ITS VERY IMPORTANT FOR ME NOW....PLEASE SUPPORT ME ....THANK YOU